A car passes a landmark on a highway traveling at a constant rate of kilometers per hour. Half an hour later, a second car passes the same landmark traveling in the same direction at kilometers per hour. How much time after the second car passes the landmark will it overtake the first car?
Answer by mananth(16949)

First car speed = x kmph
Second car speed = y kmph
..
second car passes the lad mark 1/2 hour later.
So in Half hour Car A has passes = x/2 km
..
They are travelling in same direction
so catch up speed = y-x kmph
..
Catch up distance = x/2 km
,...
Catch up time = (x/2) / (y-x) hours
=x/(2(y-x)hours